OpenStudy (anonymous):

at what value of x is the tangent line to y=x^2 + 3x +5 perpendicular to the line x-2y = 5

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

you are pretty much given the slope of your tangent

OpenStudy (solomonzelman):

\(\large\color{black}{x-2y = 5}\) \(\large\color{black}{-2y+x = 5}\) \(\large\color{black}{-2y= -x+5}\) \(\large\color{black}{y= (1/2)x+5/2}\) if a slope of a line that is perpendicular to yours is 1/2, what is the slope of your line?

OpenStudy (anonymous):

1/2

OpenStudy (solomonzelman):

no, 1/2 is a slope of a perpendilar line to yours. A perpendicular slope (mathematicians excuse me please) to slope m is, -1/m.

OpenStudy (anonymous):

I am not understanding this... if you have y= 1/2x +5/2 isn't the slope 1/2

OpenStudy (solomonzelman):

yes. but this is the rearranged version of a line that is perpendicular to your tangent that you are finding.

OpenStudy (solomonzelman):

I want you to give me a slope of a line that is perpendicular to the line with slope of 1/2.

OpenStudy (anonymous):

2?

OpenStudy (solomonzelman):

what is the sign?

OpenStudy (anonymous):

ugh... I am not sure

OpenStudy (anonymous):

positive ?

OpenStudy (solomonzelman):

-2, no? m -> -1/m 1/2 -> -2/1 -> -2

OpenStudy (anonymous):

oh I see what you mean with -1/m now , ok

OpenStudy (solomonzelman):

Yes, and now tell me what is the derivative of \(\large\color{black}{y=x^2 + 3x +5}\)

OpenStudy (anonymous):

2x +3

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

so I plug in -2 as x?

OpenStudy (solomonzelman):

the derivative of the function is the slope of it. So therefore, knowing that the slope of the tangent is -2. then you set f'(x)=-2.

OpenStudy (solomonzelman):

\(\large\color{black}{f'(x)=2x+3}\) \(\large\color{black}{-2=2x+3}\)

OpenStudy (anonymous):

and that is the equation or would you continue ?

OpenStudy (anonymous):

i mean no , I am not looking for an equation I am looking for the value of x so I continue to find x correct?

OpenStudy (solomonzelman):

no that is not the equation, you are trying to find the x-coordinate

OpenStudy (solomonzelman):

x =?

OpenStudy (anonymous):

-2/5

OpenStudy (anonymous):

no wait

OpenStudy (anonymous):

-5/2

OpenStudy (anonymous):

right?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

now find f(5/2)

OpenStudy (anonymous):

omg yay ... I get it

OpenStudy (anonymous):

ok , so now I plug that in to the y= equation right?

OpenStudy (solomonzelman):

yes you need the f(5/2) found it?

OpenStudy (solomonzelman):

I mean f(-5/2)

OpenStudy (anonymous):

15/4

OpenStudy (solomonzelman):

yes, that is what I am getting to

OpenStudy (solomonzelman):

use poitn slope formula

OpenStudy (solomonzelman):

verification. https://www.desmos.com/calculator/3nt7l5veti

OpenStudy (anonymous):

so the vlue of x is 15/4?

OpenStudy (solomonzelman):

no that is the f(-5/2)

OpenStudy (solomonzelman):

that is the y value of the point (on a function), where the x=-5/2.

OpenStudy (anonymous):

so what do I do now?

OpenStudy (solomonzelman):

you plug in your point, (-5/2 , 15/4) and slope -2, into the point slope formula.

OpenStudy (solomonzelman):

as I did in my desmos attachmentr

OpenStudy (anonymous):

yeah that is what I got

OpenStudy (solomonzelman):

yes, and you can see that it is correct. since it is tangent according to the graph. https://www.desmos.com/calculator/zxm7radxgc

OpenStudy (solomonzelman):

I just changed the color in this link

OpenStudy (anonymous):

so I am done with the problem right?

OpenStudy (solomonzelman):

questions about this problem?

OpenStudy (solomonzelman):

I got to go now, it is 10:30 pm

OpenStudy (solomonzelman):

see you!

OpenStudy (anonymous):

thank you for your help !!!